determine the amount of heat, in kJ, required to convert 12 g of solid Ca at 355 C to its melting point of 851 C
Answer #1

Sol. As Mass of Ca = 12g

Molar Mass of Ca = 40.08 g/mol

So , Moles of Ca = n = 12 / 40.08 = 0.29940  mol  

Total heat required

= Heat required to raise the temperature of solid Ca from 355°C to 851°C + Heat required to convert solid Ca to liquid Ca at 851°C

= n × Cs × deltaT + n × deltaHfus

where

Cs = Specific heat capacity of solid Ca = 0.0262 KJ/mol °C

deltaT = change in temp. from 355°C to 851°C =

(851 - 355 )°C = 496 °C   

and , deltaHfus = Enthalpy of fusion of Ca = 9.33 KJ/mol

So , Total heat required   

= 0.29940 × 0.0262 × 496 + 0.29940 × 9.33

= 6.6842 KJ